# Revisiting the relaxation of constraints in gauge theories

### v1

@@ 1 Introduction
In recent preprints [1, 2, 3, 4], an idea of relaxing the Lagrangian constraints of gauge theories has been put forward. The timeflow of these claims was as follows. The initial paper [1] dealt with the problem of time in quantum cosmology, or more precisely with the rather natural fact that its Hamiltonian vanishes on-shell. For sure, relaxing the first Friedmann equation does allow us to make the on-shell Hamiltonian non-zero. Then the next papers [2, 3] discussed relaxing the constraints in full General Relativity and in electromagnetism respectively, while the last paper in the series [4] continued considering some cosmological implications.
@@
We find it interesting to add yet another aspect to this theme. The paper [5] is written in the Hamiltonian language or, one can say, in terms of the first-order action functionals. We show that, starting from the proper Lagrangian description of gauge theories [^1] As we will see below, it is only for those of them in which the ”gauge hits twice” [6]., the relaxation happens when we take the primary constraints only, and fix the gauge by adding conditions which are of second class with them.
@@ 2 The simple case of electrodynamics in vacuum
Let us first discuss the most classical field theory, that is electrodynamics. For the sake of simplicity, we take it in vacuum. The dynamical variable is a vector field $A^{\mu}$ and the field strength is defined as
@@
where the Lagrange multiplier imposes the primary constraint of $\pi_{0}=0$. Taking the usual definition of spatial momenta, $\pi_{i}=\frac{\partial\mathcal{L}}{\partial{\dot{A}}_{i}}$, we immediately see that finally both pictures produce the same equations
@@
Actually, one of the ways to quantise a gauge model is to take the extended Hamiltonian and fix the gauge [6], i.e. to add new constraints which would be a set of second class with the real ones. One can even do so without violating the Lagrangian equations. For that we need to choose a gauge with $\tilde{\lambda}=0$. One option would be to take $A_{0}=\partial_{i}A_{i}=0$ (recall that we are in vacuum):
${\mathcal{H}}_{\mathrm{fixed}}=\frac{1}{2}\pi^{2}_{i}+\pi_{i}\partial_{i}A_{0}+\frac{1}{4}F_{ij}^{2}+\lambda\pi_{0}+\tilde{\lambda}\partial_{i}\pi_{i}+\lambda^{\prime}A_{0}+{\tilde{\lambda}}^{\prime}\partial_{i}A_{i}.$
One can check that, given suitable boundary conditions at spatial infinity, there is no gauge freedom left, for the equations imply $\lambda=\lambda^{\prime}=\bigtriangleup\tilde{\lambda}=\bigtriangleup\tilde{\lambda}^{\prime}=0$, and the usual Maxwell equations for $F_{\mu\nu}$ are there.
@@
${\mathcal{H}}_{\mathrm{relaxed}}=\frac{1}{2}\pi^{2}_{i}+\pi_{i}\partial_{i}A_{0}+\frac{1}{4}F_{ij}^{2}+\lambda\pi_{0}+\lambda^{\prime}A_{0},$ (4)
@@
Due to the last equation, the longitudinal component could then be counted as one half dynamical degree of freedom. Actually, this is nothing but the remaining gauge freedom of $\delta A_{\mu}=\partial_{\mu}\phi(x_{i})$. However, when imposed directly inside the action functional, it leads to ${\ddot{A}}_{i}=\partial_{j}F_{ji}$ being the only equation. In this case, we can only conclude that $\partial_{i}{\ddot{A}}_{i}=0$. In other words, all three modes are fully dynamical, i.e. one full new degree of freedom, only one half of which belongs to what was the gauge freedom. Formally, for the gauge invariant variables $F_{\mu\nu}$, we have got one half new degree of freedom (5) since the function $\rho(x_{i})$ acts as one new Cauchy datum.
@@ 4 The general arguments for mechanical systems
An important point about electrodynamics above is that its gauge freedom ”hits twice” [6]. The gauge transformation mixes different types of derivatives. The $A_{0}$ component lacks the time derivative in the action, while another component, $A_{i}$, lacks it own $\partial_{i}$ derivative there. Therefore, the primary constraint is not a gauge generator in itself, unless we go for the extended Hamiltonian, and generates the secondary constraint which can finally be translated into the Lagrangian one. Only their combination then serves as a generator of gauge transformations in the initial meaning [13].
@@
Let us now argue that the situation we have found in the case of electrodynamics is very general for any gauge symmetries hitting twice [^4] Somewhat artificially, one can construct systems in which it hits more than twice, for example the Lagrangian [13] of $L={\dot{x}}{\dot{z}}+yz$ where the primary constraint $p_{y}=0$ entails two secondary (or, one can say, a secondary and a tertiary) constraints $z=0$ and $p_{x}=0$, all of first class. Fixing $y=0$ inside the action removes them both. Up to a boundary term, it was a gauge symmetry of $\delta x=\epsilon(t)$, $\delta y=\ddot{\epsilon}(t)$.. For simplicity, consider a mechanical model
@@
A toy model. In a simple model [7] of $L=\frac{1}{2}({\dot{q}}+s)^{2}$ the $s$-equation ${\dot{q}}+s=0$ is indeed stronger than the $q$-equation ${\ddot{q}}+\dot{s}=0$. There is no dynamical mode, one gauge freedom, so that $q$ can be taken as pure gauge, and one fully constrained physical mode represented by the condition of $s=-\dot{q}$. In the Hamiltonian language, $H_{T}=\frac{1}{2}p_{q}^{2}-sp_{q}+\lambda p_{s},$ the primary constraint of $p_{s}=0$ entails the secondary one of $p_{q}=0$, and then Hamiltonian equations also reduce to ${\dot{q}}=-s$.
Putting a relaxation gauge of $s=0$ into the equations results in $\dot{q}=0$ which featues one half fake dynamical degree of freedom of the remnant gauge symmetry. Imposing it at a higher theoretical level does change the model. In Lagrangian formalism, it gives $L=\frac{1}{2}{\dot{q}}^{2}$ and therefore $\ddot{q}=0$. At the same time, $\tilde{\lambda}s$ added to the total Hamiltonian cancels any secondary constraint appearance and also leads to $\ddot{q}=0$. This is one degree of freedom, only one half of which is in the remnant gauge symmetry, while another half is a genuinely new content.
@@ 6 Discussion and Conclusions
All in all, what happens in the relaxation procedure is that we choose to impose a gauge fixing condition before looking for the secondary constraint. Moreover, we insist on using an incomplete gauge fixing thus producing a remnant gauge freedom and some fully new dynamics if the condition was imposed before deriving the equations of motion. Being of the lower time-derivative order, the momentum-constrained variable has naturally got a time derivative of the gauge parameter in its gauge transformation, and therefore fixing the gauge in terms of such variable is always incomplete, whatever boundary condition we might have imposed at the spatial infinity.
Gauge symmetry is an important concept in theoretical physics. Among other things, it allows us to write a nice action principle for equations of interaction fields. Therefore, removing it at the foundational level is a strange idea, to start with, let alone doing it with a gauge fixing which is necessarily incomplete even in perturbation theory. At the same time, an absolutely full gauge fixing would be difficult to naturally define in a field theory, and globally it is just impossible [21] for interesting non-Abelian theories due to Gribov ambiguity [^6] This is a topological obstruction [21]. Gribov himself found it [22] as a special case of remaining gauge freedom after imposing the Lorenz condition $\partial_{\mu}A_{a}^{\mu}=0$ in Euclidean time with boundary conditions at infinity..

### v2

@@ 1 Introduction
In recent preprints [1, 2, 3, 4], an idea of relaxing the Lagrangian constraints in gauge theories has been put forward. The timeflow of these claims was as follows. The initial paper [1] dealt with the problem of time in quantum cosmology, or more precisely with the rather natural fact that its Hamiltonian vanishes on-shell. For sure, relaxing the first Friedmann equation does allow us to make the on-shell Hamiltonian non-zero. Then the next papers [2, 3] discussed relaxing the constraints in full General Relativity and in electromagnetism respectively, while the last paper in the series [4] continued considering some cosmological implications.
@@
We find it interesting to add yet another aspect to this theme. The paper [5] is written in the Hamiltonian language or, one can say, in terms of the first-order action functionals. We show that, starting from the proper Lagrangian description of gauge theories, the relaxation happens [^1] As we will see below, it is only for those of them in which the ”gauge hits twice” [6]. when we take the primary constraints only, and fix the gauge by adding conditions which are of second class with them.
@@ 2 The simple case of electrodynamics in vacuum
Let us first discuss the most classical field theory, that of electrodynamics. For the sake of simplicity, we take it in vacuum. The dynamical variable is a vector field $A^{\mu}$ and the field strength is defined as
@@
where the Lagrange multiplier $\lambda$ imposes the primary constraint of $\pi_{0}=0$. Taking the usual definition of spatial momenta, $\pi_{i}=\frac{\partial\mathcal{L}}{\partial{\dot{A}}_{i}}$, we immediately see that finally both pictures produce the same equations
@@
Actually, one of the ways to quantise a gauge model is to take the extended Hamiltonian and fix a gauge [6], i.e. to add new constraints which would be a set of second class with the real ones. One can even do so without violating the Lagrangian equations. For that we need to choose a gauge with $\tilde{\lambda}=0$. One option would be to take $A_{0}=\partial_{i}A_{i}=0$ (recall that we are in vacuum):
${\mathcal{H}}_{\mathrm{fixed}}=\frac{1}{2}\pi^{2}_{i}+\pi_{i}\partial_{i}A_{0}+\frac{1}{4}F_{ij}^{2}+\lambda\pi_{0}+\tilde{\lambda}\partial_{i}\pi_{i}+\utilde{\lambda}A_{0}+\utilde{\tilde{\lambda}}\partial_{i}A_{i}.$
One can check that, given suitable boundary conditions at spatial infinity, there is no gauge freedom left, for the equations imply $\lambda=\utilde{\lambda}=\bigtriangleup\tilde{\lambda}=\bigtriangleup\utilde{\tilde{\lambda}}=0$, and the usual Maxwell equations for $F_{\mu\nu}$ are there.
@@
${\mathcal{H}}_{\mathrm{relaxed}}=\frac{1}{2}\pi^{2}_{i}+\pi_{i}\partial_{i}A_{0}+\frac{1}{4}F_{ij}^{2}+\lambda\pi_{0}+\utilde{\lambda}A_{0},$ (4)
@@
Due to the last equation, the longitudinal component could then be counted as one half dynamical degree of freedom. Actually, this is nothing but the remaining gauge freedom of $\delta A_{\mu}=\partial_{\mu}\phi(x_{i})$. However, when imposed directly inside the action functional, it leads to ${\ddot{A}}_{i}=\partial_{j}F_{ji}$ being the only equation. In this case, we can only conclude that $\partial_{i}{\ddot{A}}_{i}=0$. In other words, all three modes are fully dynamical, i.e. one full new degree of freedom, only one half of which belongs to what was the gauge freedom. Formally, for the gauge invariant variables $F_{\mu\nu}$, we have got one half new dynamical degree of freedom (5) since the function $\rho(x_{i})$ acts as one new Cauchy datum.
@@ 4 The general arguments for mechanical systems
An important point about electrodynamics above is that its gauge freedom ”hits twice” [6]. The gauge transformation mixes different types of derivatives. The $A_{0}$ component lacks the time derivative in the action, while another component, $A_{i}$, lacks it own $\partial_{i}$ derivative there. Therefore, the primary constraint is not a gauge generator in itself, unless we go for the extended Hamiltonian, and it generates the secondary constraint which can finally be translated into the Lagrangian one. Only their combination then serves as a generator of gauge transformations in the initial meaning [13].
@@
Let us now argue that the situation we have found in the case of electrodynamics is very general for any gauge symmetries hitting twice [^4] Somewhat artificially, one can construct systems in which it hits more than twice, for example a Lagrangian [13] of $L={\dot{x}}{\dot{z}}+yz$ where the primary constraint $p_{y}=0$ entails two secondary (or, one can say, a secondary and a tertiary) constraints $z=0$ and $p_{x}=0$, all of first class. Fixing $y=0$ inside the action removes them both. Up to a boundary term, it was a gauge symmetry of $\delta x=\epsilon(t)$, $\delta y=\ddot{\epsilon}(t)$.. For simplicity, consider a mechanical model
@@
A toy model. In a simple model [7] of $L=\frac{1}{2}({\dot{q}}+s)^{2}$ the $s$-equation ${\dot{q}}+s=0$ is indeed stronger than the $q$-equation ${\ddot{q}}+\dot{s}=0$. There is no dynamical mode, one gauge freedom, so that $q$ can be taken as pure gauge, and one fully constrained physical mode represented by the condition of $s=-\dot{q}$. In the Hamiltonian language, $H_{T}=\frac{1}{2}p_{q}^{2}-sp_{q}+\lambda p_{s},$ the primary constraint of $p_{s}=0$ entails the secondary one of $p_{q}=0$, and the Hamiltonian equations also reduce to ${\dot{q}}=-s$.
Putting a relaxation gauge of $s=0$ into the equations results in $\dot{q}=0$ which featues one half fake dynamical degree of freedom of the remnant gauge symmetry. Imposing it at a higher theoretical level does change the model. In Lagrangian formalism, it gives $L=\frac{1}{2}{\dot{q}}^{2}$ and therefore $\ddot{q}=0$. At the same time, $\utilde{\lambda}s$ added to the total Hamiltonian cancels any secondary constraint appearance and also leads to $\ddot{q}=0$. This is one degree of freedom, only one half of which is in the remnant gauge symmetry, while another half is a genuinely new content.
@@ 6 Discussion and Conclusions
All in all, what happens in the relaxation procedure is that we choose to impose a gauge fixing condition before looking for the secondary constraint. Moreover, we insist on using an incomplete gauge fixing thus producing a remnant gauge freedom, and also some fully new dynamics if the condition was imposed before deriving the equations of motion. Being of the lower time-derivative order, the momentum-constrained variable has naturally got a time derivative of the gauge parameter in its gauge transformation, and therefore fixing the gauge in terms of such variable is always incomplete, whatever boundary condition we might have imposed at the spatial infinity.
Gauge symmetry is an important concept in theoretical physics. Among other things, it allows us to write a nice action principle for equations of interaction fields. Therefore, removing it at the foundational level is a strange idea, to start with, let alone doing it with a gauge fixing which is necessarily incomplete even in perturbation theory. At the same time, an absolutely full gauge fixing would be difficult to naturally define in a field theory, and globally it is just impossible [21] for physically interesting non-Abelian theories due to Gribov ambiguity [^6] This is a topological obstruction [21]. Gribov himself found it [22] as a special case of remaining gauge freedom after imposing the Lorenz condition $\partial_{\mu}A_{a}^{\mu}=0$ in Euclidean time with boundary conditions at infinity..